OpenStudy (ryanr504):

Last Question on study guide Need Help Bad. 3root Sqrt(8(cos(4pi/5)+i sin(4pi/5))

OpenStudy (awkwardequalsme):

wolframalpha.com

OpenStudy (freckles):

does say find the cube roots of such and such?

OpenStudy (ryanr504):

No it wants me to use the de morive theorem

OpenStudy (freckles):

so it isn't cube root ? it is square root?

OpenStudy (freckles):

I thought it was cube root because of the 3 but should I ignore the 3?

OpenStudy (freckles):

I'm just asking what does 3root sqrt mean? If it doesn't mean cube root or square root, I'm totally lost.

OpenStudy (ryanr504):

No my bad it is cube im sorry im just confused

OpenStudy (freckles):

\[\text{ Let } z=r(\cos(\theta)+i \sin(\theta)) \\ \\ \text{ But } z \text{ actually can be represented in infinitely many ways } \\ \\ \text{ that is, it can be written as } z=r (\cos(\theta+2 n \pi)+i \sin(\theta+2 n \pi))\] \[\text{ Now if wanted to find the } k-\text{th root of } z\] \[\text{ there are actually } k \text{ amount of } k-\text{th roots of } z\] Then I have \[z^\frac{1}{k}=r^{\frac{1}{k}} (\cos(\frac{ \theta+ 2 n \pi}{k})+i \sin (\frac{\theta+2n \pi}{k})) \\ n=0,1,...,k-1\]

OpenStudy (freckles):

Your book should say something like this

OpenStudy (freckles):

You will wind up with 3 roots

OpenStudy (freckles):

Just use the formula in the theorem

OpenStudy (freckles):

your theta is 4pi/5

OpenStudy (freckles):

your k is 3

OpenStudy (ryanr504):

This is the formula that i have for it (cos θ + i sin θ)n = cos nθ + i sin nθ.

OpenStudy (freckles):

\[(\cos(\theta) +i \sin(\theta))^n=\cos(n \theta)+i \sin (n \theta)\] this is correct except here we have n=1/k \[(\cos(\theta)+i \sin(\theta))^\frac{1}{k}=\cos(\frac{1}{k} \theta)+i \sin(\frac{1}{k} \theta) \\ (\cos(\theta)+i \sin( \theta))^\frac{1}{k} = \cos(\frac{\theta}{k})+i \sin (\frac{\theta}{k})\]

OpenStudy (freckles):

were your k is 3 in this case

OpenStudy (ryanr504):

ok so it would be (cos(4pi/5)+i sin(4pi/4)1/3=cos(-.81/3)+i sin(.59/3)?

OpenStudy (freckles):

you will also need to use the period is 2pi to find the other 2 cube roots I'm not sure how 4pi/5 turned into -.81 and .59 at the same time :p

OpenStudy (ryanr504):

I found the exact value of cos and sin (4pi/5)

OpenStudy (freckles):

\[(\cos(\frac{4 \pi}{5})+ i \sin(\frac{4\pi}{5}))^\frac{1}{3} = \cos(\frac{1}{3}\frac{4\pi}{5})+i \sin(\frac{1}{3} \frac{4 \pi}{5})\]

OpenStudy (freckles):

I should put a multiplication sign there between the fractions \[(\cos(\frac{4 \pi}{5})+ i \sin(\frac{4\pi}{5}))^\frac{1}{3} = \cos(\frac{1}{3} \cdot \frac{4\pi}{5})+i \sin(\frac{1}{3} \cdot \frac{4 \pi}{5})\]

OpenStudy (freckles):

anyways this will be one cube root there are still two others this goes back to using the period is 2pi

OpenStudy (ryanr504):

Ok i see where i messed up

OpenStudy (freckles):

\[z=8 (\cos(\frac{4 \pi}{5}+2 n \pi)+i \sin( \frac{4\pi}{5}+2 n \pi)) \\ z^\frac{1}{3}=[8 ( \cos(\frac{4 \pi}{5}+2n \pi)+i \sin(\frac{4\pi}{5}+2 n \pi)] ^\frac{1}{3} \\ z^\frac{1}{3}=8^\frac{1}{3} (\cos(\frac{1}{3}(\frac{4 \pi}{5}+2 n \pi))+ i \sin(\frac{1}{3}(\frac{4 \pi}{5}+2 n \pi)))\]

OpenStudy (ryanr504):

Then would i multiply from there or am i still missing a step?

OpenStudy (freckles):

yes distribute the 1/3 inside both of those trig function things

OpenStudy (freckles):

and then to find the 3 cube roots just enter in n=0,1,2

OpenStudy (ryanr504):

Ok i would where would i enter n=0 in place of the z?

OpenStudy (freckles):

do you see the n?

OpenStudy (freckles):

that is where you replace n at...

OpenStudy (ryanr504):

Oh ok im sorry

OpenStudy (freckles):

\[z=8 (\cos(\frac{4 \pi}{5}+2 n \pi)+i \sin( \frac{4\pi}{5}+2 n \pi)) \\ z^\frac{1}{3}=[8 ( \cos(\frac{4 \pi}{5}+2n \pi)+i \sin(\frac{4\pi}{5}+2 n \pi)] ^\frac{1}{3} \\ z^\frac{1}{3}=8^\frac{1}{3} (\cos(\frac{1}{3}(\frac{4 \pi}{5}+2 \color{red}{(0)} \pi))+ i \sin(\frac{1}{3}(\frac{4 \pi}{5}+2 \color{red}{(0)} \pi)))\]

OpenStudy (freckles):

and you would do this for n=1 and n=2

OpenStudy (ryanr504):

i got z=20.19+6.99i for the 3 cube

OpenStudy (ryanr504):

n=2 i got 14.25+4.93i

OpenStudy (freckles):

how did you those

OpenStudy (ryanr504):

i replaced n and the multiply cos(1/3)(4pi/5+2(3)pi)

OpenStudy (freckles):

\[z_n^\frac{1}{3}=8^\frac{1}{3}( \cos(\frac{1}{3}(\frac{4 \pi}{5}+2 n \pi))+i \sin(\frac{1}{3} (\frac{4\pi}{5}+2 n \pi)) \\ z_0^\frac{1}{3}=2 (\cos(\frac{4\pi}{15})+ i \sin(\frac{4\pi}{15})) \\ z_1^\frac{1}{3}=2 ( \cos(\frac{4\pi}{15}+\frac{2 \pi}{3})+ i \sin(\frac{4\pi}{15} + \frac{ 2 \pi}{3}) ) \\ z_1^\frac{1}{3}=2( \cos(\frac{14 \pi}{15})+ i \sin(\frac{14 \pi}{15})) \] ok I will leave the last cube root to you (the simplifying of fractions part) but anyways try putting these into your calculator again

OpenStudy (ryanr504):

Ok thank you i understand this alot better now ill let you know if i need help again

OpenStudy (freckles):

k

OpenStudy (freckles):

and I see what you did also you entered in n=3 from some reason (this will be the same answer at n=0) you did cos(1/3) *(4pi/5+2*3pi) instead of cos(1/3(4pi/5+2*3pi)) the whole thing needs to be inside the cos not just the 1/3

OpenStudy (ryanr504):

OK thanks again

OpenStudy (freckles):

np good luck ryan